OpenStudy (anonymous):

verify that (cscx-cotx)^2 = 1-cosx/1+cosx

OpenStudy (anonymous):

\[\begin{align*}(\csc x-\cot x)^2&=\frac{1-\cos x}{1+\cos x}\\ &=\frac{1-\cos x}{1+\cos x}\cdot\frac{1-\cos x}{1-\cos x}\\ &=\frac{(1-\cos x)^2}{(1-\cos^2 x}\end{align*}\] What can you do with this?

OpenStudy (anonymous):

can you multiply by the conjugate

OpenStudy (anonymous):

Yeah, that's what I did. (1 - cosx) is the conjugate of (1 + cosx).